Grade 12th passMechanics

An aeroplane moving in a circular path with a speed 250 km/h.The change in velocity in half of the revolution

Saurabh sinha
8 Years agoGrade 12th pass

1 Answer

Venkat
8 Years ago
Calculate the change in velocity  as 2vsin(theta)/2
For half revolution, the theta is 180
So, change in velocity is 2vsin(180/2) = 2vsin90 = 2v = 2*250 = 500 km/hr